# Remote Sensing Spatio-Temporal Vision-Language Models: A Comprehensive Survey

### v2

@@ III-B Key technologies in RS-STVLMs
Remote sensing spatio-temporal vision-language models rely on several core technical components that collectively enable effective spatio-temporal semantic understanding for remote sensing temporal image sequences. These general technologies can be applied across different tasks. Unlike the task-specific discussions in previous sections, this section focuses on the fundamental technical blocks across various models, highlighting key methodologies used for temporal image encoding, spatio-temporal visual feature fusion, language encoding, vision-language interaction, and language generation.
@@ III-B1 Temporal Image Encoding
Temporal image encoding is critical for extracting robust spatial-temporal visual features from remote-sensing temporal image sequences. Convolutional Neural Networks (CNNs) (e.g., VGG [149], ResNet [150]), have been extensively employed as backbone architectures due to their strength in capturing local textures and fine-grained spatial details. Compared to CNNs, Transformer-based Vision models (e.g. ViT [151], Swin Transformer [152], Segformer [153], SAM [154]) have gained prominence because of their ability to capture global context via self-attention mechanisms, which is especially beneficial for modeling long-range dependencies in complex remote sensing scenes.
@@
Most approaches initialize these encoders with pre-trained weights obtained from large-scale vision datasets using either supervised pre-training (on ImageNet [155], SA-1B [154] or some remote sensing datasets) or self-supervised pre-training methods (e.g., MAE [156]). In addition, contrastive vision-language pre-training methods (e.g., CLIP [157], SigLip [158], RemoteCLIP [159], GeoRSCLIP [160] and Git-RSCLIP [161]) have also proven effective for transferring learned representations to remote sensing tasks. For instance, Liu et al. [72] compared the performance of ViT models pretrained on ImageNet and CLIP, demonstrating that vision-language pre-alignment is critical for RS-STVLMs.
@@
Beyond single-image encoders, some recent studies treat temporal image sequences as continuous videos, employing a unified video encoder to capture spatio-temporal dynamics in a holistic manner [85, 113, 89]. For example, Change3D [89] treats bi-temporal images as mini-videos consisting of two frames and integrates some learnable frames between the two images to facilitate the video encoder X3D-L [162] to perceive objects. Zhou et al. [27] proposes a single-stream extractor pre-trained on a large-scale bi-temporal remote sensing image dataset through temporal contrastive pre-training, significantly enhancing the robustness of change feature extraction.
III-B2 Spatio-Temporal Visual Feature Fusion
@@
Spatio-temporal visual feature fusion is critical for integrating visual features extracted from images acquired at different time points. The goal is to effectively capture meaningful dynamic changes while mitigating irrelevant variations caused by factors such as lighting, atmospheric conditions, or weather effects. Effective spatio-temporal fusion mechanisms ensure that RS-STVLMs capture spatial context to identify spatial changed objects and capture temporal dependencies to recognize temporal evolution patterns such as object emergence, disappearance, or expansion.
@@
Early research primarily focused on identifying and localizing changes at the spatial level. These methods often relied on direct spatial feature differencing between adjacent timestamps, assuming that feature-level differences reflect changed regions. To improve the robustness of change perception, many approaches introduced various spatio-temporal feature fusion modules. For example, Sun et al. [95] designed a diffusion-based feature extraction module to learn the change feature distribution and a frequency-guided filter module for noise suppression in high-dimensional distribution. Zhu et al. [102] integrated a bi-temporal change semantic filter into the SAM-based encoder, which combines spatial filters and channel filters to facilitate the filtration of temporal change semantic features. Yang et al. [87] proposed a time-channel-spatial attention module to dynamically weight bi-temporal features, enabling models to focus on critical change regions. However, these approaches focused more on spatial-level change perception, overlooking the necessity of explicitly modeling inter-temporal relationships.
@@
Some researchers realize that real-world remote sensing temporal scenarios involve not just identifying what has changed spatially, but also how these changes evolve over time (emergence, disappearance, or expansion). This realization has led to a focus shift towards joint spatio-temporal modeling. Recent methods have introduced temporal reasoning mechanisms to capture complex inter-temporal dependencies, enabling models to reason about dynamic change patterns across time. For instance, Li et al. [84] designed an inter-temporal cross-temporal attention module to aggregate the genuine changes and a symmetric difference transformer to leverage temporal symmetry. Liu et al. [75] proposed Spatial Difference-aware SSM (SD-SSM) and Temporal-Traversing SSM (TT-SSM) to improve joint spatio-temporal modeling. Li et al. [116] customized a change extraction module to enhance spatio-temporal relationship features between image pairs. Zou et al. [107] proposed a Frequency–Spatial–Temporal Fusion Network to disentangle structural changes and capture evolutionary patterns of geospatial changes across spatio-temporal dimensions.
@@
It is worth noting that the above-mentioned spatio-temporal visual feature fusion generally exists in methods that use multiple image encoders to process multiple images separately. This can also be called post-feature fusion. Some emerging approaches (e.g., [89, 85, 113]) explored using pre-trained video encoders, which treat multi-temporal image sequences as continuous temporal streams. Unlike previous approaches that require handcrafted spatio-temporal fusion modules, video encoders naturally capture both spatial and temporal patterns through end-to-end feature extraction, without requiring explicit spatio-temporal feature fusion strategies.
@@ III-B3 Language Encoding
The language encoder transforms textual inputs, such as user queries or descriptive questions, into semantic embeddings that the model can effectively process. This step is essential for ensuring the model can accurately interpret and respond to user instructions. In the context of remote sensing spatio-temporal VLMs, most methods rely on off-the-shelf pre-trained models followed by task-specific fine-tuning. Pre-trained language models, such as BERT [117] and T5 [141], are commonly used due to their proven ability to capture nuanced language representations. These models leverage extensive training on large-scale corpora, enabling them to encode complex, context-rich sentences into fixed-length vectors that encapsulate essential semantic information. In addition, it is also a good choice to use the text encoder of CLIP pretrain on the remote sensing vision-language dataset, such as RemoteCLIP [159], GeoRSCLIP [160] and Git-RSCLIP [161].
@@ III-B4 Vision-Language Interaction
A core challenge in spatio-temporal vision-language models is the effective fusion of visual and textual modalities. The vision-language interaction module is responsible for bridging the semantic information derived from multi-spatio-temporal visual features with that from textual inputs, enabling the model to perform tasks such as change question answering and change grounding.
Various strategies have been proposed to achieve this interaction. Some approaches adopt an early interaction paradigm, where visual features and text embeddings from the tokenizer are concatenated and fed into a unified Transformer to perform multimodal interaction. This strategy, commonly used in recent multimodal large language models (e.g., TeoChat [68], Falcon [127]), enables direct interactions between visual tokens and language tokens, facilitating joint representation learning.
Others use late interaction techniques. Specifically, the text is first processed by a pre-trained language encoder and then performs complex interactions with vision features, such as various attention mechanisms. This facilitates fine-grained interactions between visual and textual information and allows the model to refine visual features based on contextual cues from the text, thereby enhancing the model’s ability to focus on relevant change regions. For instance, Li et al. [29] used language to guide the extraction of multi-scale features and proposed a multi-stage reasoning module that facilitates fine-grained cross-modal information interaction.
@@ III-B5 Language Generation
Since it is difficult to train a language decoder from scratch on a small dataset, some recent methods have explored the use of pre-trained large language models as language decoders, such as [104, 72, 88, 127, 68]. These methods use efficient fine-tuning strategies to take advantage of the powerful text generation capabilities of large language models. The integration of LLMs into RS-STVLMs has significantly expanded the capabilities of spatio-temporal vision-language understanding. We will introduce this in detail in the section IV.
@@ IV Large Language Models Meet Temporal Images
Recent years have witnessed remarkable advances in large language models (LLMs), whose impressive abilities in text generation and comprehension have revolutionized numerous domains [164, 165, 166]. In the context of remote sensing spatio-temporal vision-language understanding, the incorporation of LLMs presents novel opportunities by integrating visual features with sophisticated language representations. This section reviews the evolution of LLMs, parameter-efficient tuning techniques, LLM-driven task-specific VLMs, unified spatio-temporal vision-language foundation models, and LLM-driven remote sensing vision-language agents.
@@ IV-A Evolution of Large Language Models
Built on the advanced Transformer architecture with scaling law[167], LLMs leverage vast amounts of corpora and computational resources to develop robust language understanding and generation capabilities. LLMs have proven capable of excelling in specific downstream tasks through fine-tuning or even performing tasks in zero-shot and few-shot settings [168, 169, 170]. From the perspective of model architecture, LLMs have evolved rapidly from early encoder-only architectures (e.g., BERT [117]) and encoder-decoder (e.g., T5 [141]) to decoder-only models (e.g., GPT series, LLaMA [171], Gemini [172], Phi [173]). Models such as ChatGPT and GPT-4 have solidified the autoregressive, Decoder-only approach as the dominant paradigm, and recent developments (e.g., GPT-4, Qwen-2.5 VL [142, 174], LLaMA-3 [175]) are increasingly extending into multimodal domains.
These models undergo a multi-phase training procedure: (1) Self-supervised pretraining (e.g., Masked Language Modeling and Autoregressive Language Modeling [176]) on large-scale unlabeled corpora to acquire general language knowledge and understanding capabilities. (2) Instruction fine-tuning using supervised datasets to improve its performance on the specific task [177]. (3) alignment with human preferences through Reinforcement Learning with Human Feedback (RLHF) [178].
@@ IV-B1 Prompt Tuning
: The method introduces trainable prompt embeddings or prefixes into the input sequence of LLMs [169], remodelling downstream tasks into tasks familiar to pre-trained LLMs to fully exploit their potential. In other words, the aim is to adapt the task to LLMs rather than adapt LLMs to the specific task.
@@ IV-B2 Adapter-Series Methods
: These methods insert additional lightweight learnable modules into specific Transformer layers [179]. Some representative methods have been developed, such as SparseAdapter [180], Adapters [181], AdaMix [182] and LeTS [183].
@@ IV-B3 LoRA (Low-Rank Adaptation)
: Instead of directly updating the weights of the pre-trained LLMs, LoRA [184] decomposes the weight update into low-rank matrices, drastically reducing the number of trainable parameters. The updated weight matrix $W_{\text{new}}$ can be expressed as:
@@
These techniques can achieve efficient adaptation of the model without the need for significant additional computational overhead. It is particularly attractive for exploring LLM-based remote sensing vision-language modeling in large-scale spatio-temporal remote sensing scenarios. Several surveys have provided a comprehensive review of efficient fine-tuning methods, such as [168, 163, 185]
Figure 10: A representative approach for LLM-driven task-specific spatio-temporal VLMs. PromptCC  [72] first introduces LLMs into the change captioning task.
IV-C LLM-driven Task-Specific Spatio-Temporal VLMs
Recent studies have begun to incorporate large language models (LLMs) into task-specific architectures to enhance performance in specific remote sensing spatio-temporal vision-language understanding tasks, such as change captioning and change question answering.

### v3

@@ III-B Key Technologies in RS-STVLMs
RS-STVLMs rely on several core technical components that collectively enable effective spatiotemporal understanding for remote sensing temporal images. These general technologies can be applied across different tasks. Unlike the task-specific discussions in previous sections, this section focuses on the fundamental technical blocks across various models, highlighting key methodologies used for temporal image encoding, spatiotemporal visual feature fusion, language encoding, vision-language interaction, and language generation.
@@ III-B1 Temporal Image Encoding
Temporal image encoding plays a crucial role in extracting robust spatiotemporal visual features from multi-temporal remote sensing image sequences. Various encoding strategies have been employed, each with distinct characteristics in capturing spatial and temporal information for RS-STVLMs.
Encoder Architecture. Convolutional Neural Networks (CNNs), such as VGG [192] and ResNet [193], have been widely used due to their ability to capture local textures and fine-grained spatial details. However, their inherent locality and limited receptive field hinder their effectiveness in modeling long-range spatial dependencies and temporal dynamics, making them less suitable for complex temporal reasoning tasks. Transformer models like ViT [194], Swin Transformer [195], Segformer [196], and SAM [197] leverage self-attention mechanisms to capture global context and long-range dependencies. This makes them more adaptable to heterogeneous and large-scale remote sensing scenes. Their flexibility and scalability also enable integration with temporal modeling modules, which enhances their applicability in RS-STVLMs. However, they often require large-scale pre-training and significant computational resources.
@@
Pre-trained Image Encoders. Most RS-STVLMs leverage pre-trained image encoders via supervised (e.g., ImageNet [198], SA-1B [197], some remote sensing datasets) or self-supervised approaches (e.g., MAE [199]). In addition, contrastive vision-language pre-training methods (e.g., CLIP [200], SigLip [201], RemoteCLIP [202], GeoRSCLIP [203] and Git-RSCLIP [204]) have also proven effective for transferring learned representations to remote sensing tasks. Liu et al. [98] show that CLIP-pretrained ViT outperforms traditional ImageNet-pretrained counterparts in change captioning tasks, demonstrating that vision-language pre-alignment is critical for RS-STVLMs.
@@
Video-based Encoders. To encode temporal sequences more holistically, beyond image-based encoders, some recent works treat multi-temporal image sequences as video input and adopt video-specific encoders to capture spatiotemporal dynamics [111, 139, 115]. For example, Change3D [115] treats bi-temporal images as mini-videos and integrates some learnable frames between the two images to facilitate the video encoder X3D-L [205] to perceive frame-wise changes. Zhou et al. [29] pre-trained a single-stream extractor on a large-scale bi-temporal image dataset through temporal contrastive pre-training, significantly enhancing temporal change sensitivity. Compared with traditional image-encoder based methods that require explicit temporal fusion strategies (e.g., feature concatenation, temporal attention), video encoders inherently learn temporal dependencies in an end-to-end manner, reducing the need for handcrafted fusion designs. However, they are computationally intensive and are harder to pretrain due to limited video-like remote sensing datasets.
III-B2 SpatioTemporal Visual Feature Fusion
@@
Spatiotemporal visual feature fusion is critical for integrating visual features extracted from images acquired at different time points. The goal is to effectively capture meaningful dynamic changes while mitigating irrelevant variations caused by factors such as lighting, atmospheric conditions, or weather effects. Effective spatiotemporal fusion mechanisms ensure that RS-STVLMs capture spatial context to identify spatial changed objects and capture temporal dependencies to recognize temporal evolution patterns such as object emergence, disappearance, or expansion.
@@
Early research primarily focused on identifying and localizing changes at the spatial level. These methods often relied on direct spatial feature differencing between adjacent timestamps, assuming that feature-level differences reflect changed regions. To improve the robustness of change perception, many approaches introduced various spatiotemporal feature fusion modules. For example, Sun et al. [121] designed a diffusion-based feature extraction module to learn the change feature distribution and a frequency-guided filter module for noise suppression in high-dimensional distribution. Zhu et al. [128] integrated a bi-temporal change semantic filter into the SAM-based encoder, which combines spatial filters and channel filters to facilitate the filtration of temporal change semantic features. Yang et al. [113] proposed a time-channel-spatial attention module to dynamically weight bi-temporal features, enabling models to focus on critical change regions. However, these approaches focused more on spatial-level change perception, overlooking the necessity of explicitly modeling inter-temporal relationships.
@@
Some researchers realize that real-world remote sensing temporal scenarios involve not just identifying what has changed spatially, but also how these changes evolve over time (emergence, disappearance, or expansion) [101, 110, 133, 206]. This realization has led to a focus shift towards joint spatiotemporal modeling. Recent methods have introduced temporal reasoning mechanisms to capture complex inter-temporal dependencies, enabling models to reason about dynamic change patterns across time. For instance, Li et al. [110] designed an inter-temporal cross-temporal attention module to aggregate the genuine changes and a symmetric difference transformer to leverage temporal symmetry. Liu et al. [101] proposed Spatial Difference-aware SSM (SD-SSM) and Temporal-Traversing SSM (TT-SSM) to improve joint spatiotemporal modeling. Li et al. [142] customized a change extraction module to enhance spatiotemporal relationship features between image pairs. Zou et al. [133] proposed a Frequency-Spatial-Temporal Fusion Network to disentangle structural changes and capture evolutionary patterns of geospatial changes across spatiotemporal dimensions.
@@
It is worth noting that the above-mentioned spatiotemporal visual feature fusion generally exists in methods that use multiple image encoders to process multiple images separately. This can also be called post-feature fusion. Some emerging approaches (e.g., [115, 111, 139]) explored using pre-trained video encoders, which treat multi-temporal image sequences as continuous temporal streams. Unlike previous approaches that require handcrafted spatiotemporal fusion modules, video encoders naturally capture both spatial and temporal patterns through end-to-end feature extraction, without requiring explicit spatiotemporal feature fusion strategies.
@@ III-B3 Language Encoding
The language encoder transforms textual inputs, such as user queries or descriptions, into semantic embeddings that RS-STVLMs can effectively interpret. This component is essential for enabling accurate language understanding and facilitating cross-modal interaction.
Many existing RS-STVLMs employ pre-trained language models such as BERT [144] or T5 [207], followed by task-specific fine-tuning. These models, trained on large-scale text corpora, exhibit strong capabilities in understanding contextual semantics and performing complex language reasoning. With the rise of LLMs, more powerful linguistic representations are now accessible. However, LLMs are typically trained only on textual data and lack inherent multi-modal alignment, making it challenging to bridge the semantic gap between language and vision modalities.
To address this, an alternative line of work leverages CLIP-like text encoders trained jointly with image encoders using contrastive learning. Remote sensing variants such as RemoteCLIP [202], GeoRSCLIP [203], and Git-RSCLIP [204] align textual and remote sensing visual representations in a shared embedding space. For instance, Git-RSCLIP [204] leverages over 10 million paired samples and 5.5 million words during pre-training, achieving SOTA performance among CLIP-like models in the remote sensing community. Benefiting from the vision-language pre-alignment, these CLIP-like models excel in vision-language tasks and are more data-efficient in downstream applications. However, such models are relatively limited in handling long-form textual understanding and complex reasoning compared to LLMs.
In summary, both LLMs and CLIP-like encoders offer complementary strengths. LLMs excel at language understanding and complex reasoning, but often struggle with cross-modal alignment. In contrast, CLIP-like models are optimized for vision-language alignment, yet they are relatively limited in handling linguistically complex inputs. Future research could benefit from hybrid approaches that integrate LLMs with CLIP-like encoders.
@@ III-B4 Vision-Language Interaction
A key component in RS-STVLMs is the interaction mechanism between multi-temporal visual features and textual inputs, which fundamentally determines the model’s ability to understand and reason across modalities. Existing vision-language interaction strategies can be broadly categorized into two paradigms: early interaction and late interaction.
The early interaction paradigm refers to directly concatenating visual features and language tokens (typically from embedding layers) and feeding them into a unified Transformer (e.g., a pretrained LLM) for joint modeling and implicit interaction. This approach enables dense token-level interactions from the outset, facilitating shared multimodal representation learning. It is widely adopted in recent multimodal large language models, such as TeoChat [94], RSUniVLM [140], EarthDial [141], and DVLChat [143]. However, early fusion can introduce semantic interference when modality-specific features are poorly aligned, and it often incurs high computational costs due to the extended token sequences.
In contrast, the late interaction paradigm first encodes each modality independently using domain-specific encoders and then performs cross-modal fusion through dedicated interaction modules (e.g., cross-attention layers). This design enables modality-specialized representation learning and more flexible, controllable interaction strategies. Consequently, the design of the interaction module becomes especially critical. Recent works have proposed increasingly sophisticated cross-modal mechanisms to achieve fine-grained interaction between visual and textual modalities. For instance, Li et al. [31] used language to guide the extraction of multi-scale features and proposed a multi-stage reasoning module that leverages text to focus on relevant change regions. Huang et al. [208] introduced a text-guided interaction module that generates context-rich representations to guide the visual change perception process. Chen et al. [209] designed a cascaded second-order prompter that decomposes textual embeddings into complementary semantic subspaces to perform coarse-to-fine visual reasoning. Dong et al. [210] developed a vision-language-driven decoder that leverages contextual cues from image-text encoding to refine visual features during change decoding and strengthen the semantic relations among image-text features.
@@ III-B5 Language Generation
Since it is difficult to train a language decoder from scratch on a small dataset, some recent methods have explored the use of pre-trained LLMs as language decoders, such as [130, 98, 114, 154, 94]. These methods use efficient fine-tuning strategies to take advantage of the powerful text generation capabilities of LLMs. The integration of LLMs into RS-STVLMs has significantly expanded the capabilities of spatiotemporal vision-language understanding. We will introduce this in detail in the section IV.
@@ IV Large Language Models Meet Temporal Images
Recent years have witnessed remarkable advances in large language models (LLMs), whose impressive abilities in text generation and comprehension have revolutionized numerous domains [212, 213, 214]. In the context of remote sensing spatiotemporal vision-language understanding, the incorporation of LLMs presents novel opportunities by integrating visual features with sophisticated language representations. This section reviews the evolution of LLMs, parameter-efficient tuning techniques, LLM-driven task-specific VLMs, unified spatiotemporal vision-language foundation models, and LLM-driven remote sensing vision-language agents.
@@ IV-A Evolution of Large Language Models
Built on the advanced Transformer architecture with scaling law[215], LLMs leverage vast amounts of corpora and computational resources to develop robust language understanding and generation capabilities. LLMs have proven capable of excelling in specific downstream tasks through fine-tuning or even performing tasks in zero-shot and few-shot settings [216, 217, 218]. From the perspective of model architecture, early models utilized encoder-only architectures (e.g., BERT [144]) and encoder-decoder models (e.g., T5 [207]), while recent models like the GPT series, LLaMA [219], and Gemini [220] have adopted the decoder-only approach. The success of models like ChatGPT [221] and GPT-4 has solidified this paradigm, and the latest models (e.g., GPT-4, Qwen-2.5 VL [222], LLaMA-3 [223]) are now increasingly venturing into multimodal domains, incorporating both language and vision for complex reasoning tasks.
The typical training procedure for these models is multi-phased: 1) Self-supervised pretraining on large-scale unlabeled corpora, involving tasks like Masked Language Modeling and Autoregressive Language Modeling [224], to build general language understanding. 2) Instruction fine-tuning using supervised datasets, which helps improve performance on specific tasks [225]. 3) Alignment with human preferences through Reinforcement Learning with Human Feedback (RLHF) [226], further refining the model’s performance for real-world applications.
Recently, multimodal complex reasoning has emerged as a promising research direction, progressing rapidly with the advent of large multimodal reasoning models such as OpenAI’s O1/O3 and DeepSeek-R1 [227]. Reasoning models have evolved through three key stages. Early multimodal reasoning models typically used modular designs that decomposed the reasoning process into components like representation, alignment, fusion, and reasoning. For instance, Hudson et al. [228] applied memory-augmented control units for iterative reasoning. With the rise of large-scale multimodal pretraining, reasoning capabilities significantly improved, although these models still depended on surface-level pattern matching and lacked advanced capabilities for hypothesis generation and multi-step logical reasoning. For instance, IPVR [229] introduced the ”see-think-confirm” framework, guiding LLMs through visual grounding and rationale verification. Recently, there has been a shift towards deeper, long-term reasoning abilities, especially following the success of DeepSeek-R1. The GRPO algorithm of DeepSeek-R1 has become widely adopted in multimodal models, such as VLM-R1 [230], Visual-RFT [231], and Seg-Zero [232], to enhance visual capabilities like localization, detection, and classification. This reinforcement learning approach has proven effective in improving visual reasoning. Further exploration on multimodal reasoning can be found in recent surveys [233, 234, 235].
@@ IV-B1 Prompt Tuning
The method introduces trainable prompt embeddings or prefixes into the input sequence of LLMs [217], remodelling downstream tasks into tasks familiar to pre-trained LLMs to fully exploit their potential. In other words, the aim is to adapt the task to LLMs rather than adapt LLMs to the specific task.
@@ IV-B2 Adapter-Series Methods
These methods insert additional lightweight learnable modules into specific Transformer layers [236]. Some representative methods have been developed, such as SparseAdapter [237], Adapters [238], AdaMix [239] and LeTS [240].
@@ IV-B3 LoRA (Low-Rank Adaptation)
Instead of directly updating the weights of the pre-trained LLMs, LoRA [241] decomposes the weight update into low-rank matrices, drastically reducing the number of trainable parameters. The updated weight matrix $W_{\text{new}}$ can be expressed as:
@@
These techniques can achieve efficient adaptation of the model without the need for significant additional computational overhead. It is particularly attractive for exploring LLM-based remote sensing vision-language modeling in large-scale spatiotemporal remote sensing scenarios. Several surveys have provided a comprehensive review of efficient fine-tuning methods, such as [216, 211, 242]
IV-C LLM-driven Task-Specific SpatioTemporal VLMs
Recent studies have begun to incorporate LLMs into task-specific architectures to enhance performance in remote sensing spatiotemporal vision-language understanding tasks, such as change captioning and change question answering.